OpenStudy (anonymous):

Choose the graph below that correctly represents the equation 3x - 4y = 12.

OpenStudy (anonymous):

Can you get the slope on this one?

OpenStudy (anonymous):

And/or the y-intercept?

OpenStudy (anonymous):

The slope is 3/4 and the y intercept is -3

OpenStudy (anonymous):

You'de probably be able to find it easier if you switched the equation to slope intercept form, so 3x-4y=12 3x-4y=12 -3x -3x -4y=12-3x divide each side by -4 y=3/4x-3

OpenStudy (anonymous):

Then plug in the equation into your calculator and find the same graph :)

OpenStudy (anonymous):

Your slope and y-int are correct. Can you visualize that graph now? Goes through (0,-3) and goes up from left to right.

OpenStudy (anonymous):

I got it than you guys